OpenStudy (anonymous):

8x2-15=1 3+c=12 Are these true false or open sentences. I am trying to help my 7th grader

OpenStudy (mathteacher1729):

Is your first equation " eight times two - 15 = 1" ? or "eight x squared - 15 = 1" ?

OpenStudy (anonymous):

The first one is true, the second one c=9

OpenStudy (anonymous):

3+c=12 c=9

OpenStudy (anonymous):

best response!?!

OpenStudy (anonymous):

how is the first one true?

OpenStudy (anonymous):

8*2=16, add the 15 to the opposite side you get 16=16

OpenStudy (anonymous):

? that's not right i'm sorry

OpenStudy (anonymous):

YES IT IS!!!

OpenStudy (anonymous):

What is the second problem? Thank u for your help

OpenStudy (anonymous):

\[8x^2-15=1 \] \[8x^2=16\] \[x^2=2\] \[x=\sqrt{2}\]

OpenStudy (anonymous):

I could be wrong.

OpenStudy (anonymous):

Subtract 3 from both sides and you get c=9

OpenStudy (anonymous):

It is 7th grade math they dont do that

OpenStudy (anonymous):

um yes they do. clearly they are taking algebra so they must know that.

OpenStudy (anonymous):

Umm... Clearly you dont know what you are talking about

OpenStudy (anonymous):

And you completely ignored the 8X2 that could be a variable not really a multiplication sign and I do know what I'm talking about seeing as i am IN IB MATH so please don't underestimate my intelligence

OpenStudy (anonymous):

@Mmcs511 Can you please verify the question?

OpenStudy (anonymous):

It is clearly multiplication, and just because you are in that math doesn't mean you know what you are talking about...geez

OpenStudy (anonymous):

I think its 8 times 2 :)

OpenStudy (anonymous):

If they way you said is right then I give it to you. But we don't know if it's a variable or not.

OpenStudy (anonymous):

Okay then @marktj you were right.

OpenStudy (anonymous):

Thanks for mentioning me!!

OpenStudy (anonymous):

And I'm not telling you that you don't know what you're talking about. I was just saying that I thought you were wrong.

OpenStudy (anonymous):

Well im not

OpenStudy (anonymous):

Okay? And I recognized that.

OpenStudy (anonymous):

That im right?

OpenStudy (anonymous):

Yes. I recognized that you were right.

OpenStudy (anonymous):

Thank you soo much. It means a lot that i know we were able to put our egos aside and resolve this.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

No problem (: @marktj

OpenStudy (anonymous):

Thanks @jordyn77

OpenStudy (anonymous):

What Is the answer to the second proble

OpenStudy (mathteacher1729):

Since @Mmcs511 hasn't yet replied to my original question, I don't think we'll ever know 100% what exactly the first problem was asking. I *assume* that since this is 7th grade math asking about whether or not an equation is true/false, the first one means "eight TIMES 2" 8*2 -15 = 1 16 -15 = 1 1 = 1 TRUE The second equation comes with no other info. It didn't ASK to solve for c... it just said "is the equation true?" Without a value of c given, we can't tell if it's true or false. That's my two cents on it. :-p

OpenStudy (anonymous):

Dibs on best response.. LOL

OpenStudy (anonymous):

Thank you